# Entanglement measures of Majorana bound states

### v1

@@ V Conclusions
In conclusion, we have investigated the entanglement properties of Majorana bound states emerging in a topological superconductor that is coupled to two quantum dots. To contrast the impact of Majorana bound states, we have compared this Majorana system with an equivalent where the quantum dots are coupled via a normal fermion. In order to characterize entanglement in both systems, we have employed the concurrence and discord, quantum correlations that we also complemented by studying the entanglement dynamics and return probability. We found that zero-energy Majorana bound states can transform initially maximally entangled quantum dots into a classical state, while maximally entangled states are notably achieved at finite Majorana energies, thus highlighting the impact of Majorana nonlocality. Remarkably, we have shown that the maximal entanglement between quantum dots at the beginning of the evolution can be completely shift to a state between Majorana bound states and one of the quantum dots.
@@
Furthermore, we found entanglement and quantum correlations can be generated from an initially separable state in the Majorana system, with discord reliably quantifying quantum correlations for zero-energy Majorana bound states. We also showed that entanglement generation in the Majorana system is very sensitive to the nonlocality of the Majorana bound states, which, upon an appropriate control, can even induce maximally entangled states in the long time dynamics. In all the studied cases, we found that the Majorana system exhibits distinct entanglement features than the normal fermion system, which occurs because spatial nonlocality is intrinsic to Majorana bound states but absent in regular fermions. Our findings can therefore serve as an alternative way to characterize Majorana bound states and exploit their potential for quantum information.
@@ A.1 Majorana system
Using all the expressions from Eq. (32-34), the conditional entropy $C_{\theta,\phi}(\rho_{A|B_{\tilde{k}}})$ can be written by the weighted sum of marginal conditional entropies as

### v2

@@ V Concluding remarks
In conclusion, we have investigated the entanglement properties of Majorana bound states emerging in a topological superconductor that is coupled to two quantum dots. To contrast the impact of Majorana bound states, we have compared this Majorana system with an equivalent where the quantum dots are coupled via a normal fermion. In order to characterize entanglement in both systems, we have employed concurrence and discord. We also complemented this by studying the entanglement dynamics and return probability. We found that zero-energy Majorana bound states can transform initially maximally entangled quantum dots into a classical state, while maximally entangled states are notably achieved at finite Majorana energies, thus highlighting the impact of Majorana nonlocality. Remarkably, we have shown that the maximal entanglement between quantum dots at the beginning of the evolution can be completely shifted to a state between Majorana bound states and one of the quantum dots.
@@
Furthermore, we found that entanglement and quantum correlations can be generated from an initially separable state in the Majorana system, with discord reliably quantifying quantum correlations for zero-energy Majorana bound states. We also showed that entanglement generation in the Majorana system is very sensitive to the nonlocality of the Majorana bound states, which, upon an appropriate control, can even induce maximally entangled states in the long-time dynamics. The inherent Majorana nonlocality originates distinct entanglement features in the Majorana and normal fermion systems, thus highlighting the potential use of quantum correlations for distinguishing zero-energy states in Majorana devices and their utility for quantum information. Our results can be in principle measured by reconstructing the density matrix via quantum state tomography [68, 69], thus adding an alternative way for detecting Majorana states as well as offering another approach for generating maximally entanglement states beyond photon- [63, 64, 65] and cavity-mediated systems [66, 67].
It is worth noting that our results assume that the parity of the system is fixed during the time evolution, being either even or odd. Under realistic conditions, however, this situation might be difficult to achieve because nonequilibrium quasiparticles or equilibrium nontopological subgap states can poison the computational subspace and lead to parity flips, see e.g., Refs. [95, 96, 97, 98, 99]. While quasiparticle poisoning is challenging to avoid in realistic setups, it has been shown that nonequilibrium quasiparticle poisoning can be reduced with Majorana zero modes because they act as robust quasiparticle traps, giving rise to decoherence times longer than the time needed for qubit operations in moderately sized topological superconductors [99]. Moreover, the detrimental effect of quasiparticle poisoning due to nontopological subgap states, very likely due to disorder-induced states which coexist with Majorana states [27], has been shown to be negligible at weak disorder [99]. Taking these ideas into account and under similar considerations, we expect that quasiparticle poisoning can be mitigated in the systems we study. We, however, anticipate that variations of the system’s parity might induce a change in the periodicity of the correlations dynamics, which can serve as a signature of quasiparticle poisoning in our case.
@@ A.1 Majorana system
Using all the expressions from Eq. (33-35), the conditional entropy $C_{\theta,\phi}(\rho_{A|B_{\tilde{k}}})$ can be written by the weighted sum of marginal conditional entropies as
@@ A.2 Normal fermion system
Appendix B Entanglement measures for the odd sector
The dynamics of the Hamiltonian in Eq. (1) can be explored in the odd sector of excitation, following the step-by-step calculation of the even sector in Sec. II. In this Appendix, we provide an explicit calculation of the odd sector Hamiltonian and its state dynamics to highlight the differences observed in this sector. The Hamiltonian matrix, $H^{o}$, in the odd (o) sector basis states, |$\ket{001}$, |$\ket{010}$, |$\ket{100}$, and |$\ket{111}$ can be written as
$H^{o}=\begin{pmatrix}\omega&\lambda&1&0\\ \lambda&-\omega&0&1\\ 1&0&-\omega&\lambda\\ 0&1&\lambda&\omega\end{pmatrix}\,,$ (39)
Figure 6: Odd sector time evolution of concurrence, discord, return probability, and entanglement dynamics for an initially maximally entangled state of QDs in the MS. (a, b) Dynamics for $\lambda=1$ and $\lambda=-1$ at $\omega=0$. (c, d) Dynamics for $\lambda=0.7$ and $\lambda=-0.7$ at $\omega=0$. Negative values of $\lambda$ and $\omega$ can invert the dynamics, switching between odd and even sectors.
where the parameters $\lambda$ is the ration of couplings of MBSs and QDs and $\omega$ is the energy associated with the overlap of the MBSs wave functions. The matrix elements equal to 1 correspond to $\lambda_{1}=1$ (see Fig. 1). This Hamiltonian matrix is similar in form to the one defined in the even sector in Eq. (2), with the key difference being the sign of the coupling energy $\omega$ and $\lambda_{1}$. However, this sign change does not affect the eigenvalues, which are still given by the same equation as in Eq. (3). Therefore, the energy eigenvalues for the odd sector Hamiltonian can be written as
$E^{o}_{i}=\pm\sqrt{(\lambda\pm 1)^{2}+\omega^{2}}\,,$ (40)
where energy eigenvalues appear in pairs as $E^{o}_{1(2)}=\mp\Delta$ and $E^{o}_{3(4)}=\mp\Delta_{1}$, with $\Delta=\sqrt{(\lambda-1)^{2}+\omega^{2}}$ and $\Delta_{1}=\sqrt{(\lambda+1)^{2}+\omega^{2}}$. However, the eigenvectors differ from Eq. (4) in terms of local phases of different configurations. The eigenvectors |$\ket{E^{o}_{j}}$ for the paired eigenvalues in Eq. (40) can be written as un-normalized column vectors in the order of its corresponding eigenvalues as
$EV^{o}=(\ket{E^{o}_{1}}|\ket{E^{o}_{2}}|\ket{E^{o}_{3}}|\ket{E^{o}_{4}})\,,$ (41)
where |$\ket{E^{o}_{1}}=(-1,[\omega+\Delta]/[\lambda-1],-[\omega+\Delta]/[\lambda-1],1)^{T}$ and T is the transpose operation. The eigenvector |$\ket{E^{o}_{2}}=\ket{E^{o}_{1}(\Delta\rightarrow-\Delta)}$, and |$\ket{E^{o}_{3}}=(1,-[\omega+\Delta_{1}]/[\lambda+1],-[\omega+\Delta_{1}]/[\lambda+1],1)^{T}$, while |$\ket{E^{o}_{4}}=\ket{E^{o}_{3}(\Delta_{1}\rightarrow-\Delta_{1})}$. Similar to the even sector, all eigenvalues in Eq. (40) depend on $\lambda$ and $\omega$, while the corresponding eigenstates in Eq. (41) include contributions from all configurations of the odd states listed in Eq. (39). The redistribution of the local phases within the eigenstates of the odd sector differs from that in the even sector, as shown in Eq. (4). This difference affects the dynamics of entanglement measures and other dynamical quantities. Therefore, we provide calculations of the dynamical state functions in both general and specific cases to emphasize the similarities and differences in the dynamics between the two sectors.
B.1 Evolution of a maximally entangled state
In this section, we consider an initial state of maximally entangled QDs in the odd sector. For this purpose, we choose the state in the form of
$|\psi^{o}(0)\rangle=\frac{|001\rangle+|111\rangle}{\sqrt{2}}\,.$ (42)
In this initial state, the two QDs share the same form of initial entanglement as considered in Eq. (17). To preserve the odd parity in the system, the nonlocal fermion formed from the MBSs is in the state |$\ket{1}$. Subsequently, we allow the system to evolve with arbitrary values of the Hamiltonian parameters to compute all dynamical quantities. For this initial state, the system’s dynamics can be expressed as
|$\begin{split}\ket{\psi^{o}(t)}&=e^{-iE^{o}_{3}t}\ket{E^{o}_{3}}\langle E^{o}_{3}|\psi^{o}(0)\rangle\\ &+e^{-iE^{o}_{4}t}\ket{E^{o}_{4}}\langle E^{o}_{4}|\psi^{o}(0)\rangle\,,\end{split}$ (43)
where $|\psi^{o}(0)\rangle$ is the initial state defined in Eq. (42), and $|E^{o}_{3(4)}\rangle$ are given in Eq. (41). In this dynamics, the system accesses only two eigenstates, $|E^{o}_{3(4)}\rangle$, as the other two states are orthogonal to the initial state. The expression in Eq. (43) can be simplified in a similar form to the one given in Eq. (19). We can write
[…]
where $\chi^{\prime}$ and $\eta^{\prime}$ are time-dependent and given by
[…]
The expression in Eq. (44) describes the time-evolved state of the system in the odd sector and has a similar form to that in Eq. (19) for the even sector. However, the expressions for $\eta^{\prime}$ and $\chi^{\prime}$ differ from those in the even sector. These coefficients involve the eigenvalue $\Delta_{1}$ and the term $\lambda+1$, rather than $\Delta$ and $\lambda-1$ as in Eq. (20) for the even sector. We present the results in Fig. 6 for completely localized MBSs ($\omega=0$), considering both symmetric and asymmetric couplings to the QDs. In Fig. 6(a), all quantities exhibit periodic behavior except for the entanglement dynamics ($E_{d}$), which remains constant throughout. At $\omega=0$, $E_{d}$ can be expressed as $E_{d}=|e^{i(\Delta_{1}t)}|^{2}/4$, which has a constant magnitude of 0.25. In contrast, concurrence and discord show the system transitioning from maximally entangled QDs ($C=D=1$) to separable states ($C=D=0$). The return probability is also periodic, with a periodicity twice that of concurrence. This can be explained by the simple expressions $C=\cos(4t)$ and $R_{p}=\cos^{2}(2t)$. The zeros of discord can be analyzed in a manner similar to the approach used in the even sector. Furthermore, in the case of asymmetric couplings of MBSs to QDs, we present all quantities in Fig. 6(c). The dynamics are similar to those observed for symmetric coupling ($\lambda=1$), except that the periodicity changes due to a shift in the contributing eigenenergy values of the system’s Hamiltonian.
B.2 Comparing the dynamics of the odd sector correlations with their even sector counterparts
In Fig. 2(a), we have shown that the even sector dynamics of the system show no evolution of entanglement measures and the state dynamical functions at $\omega=0$ and $\lambda=1$. This feature can be retrieved in the odd sector as well by tuning the parameters carefully. The equivalent dynamics can be obtained by applying the transformations $\omega\rightarrow-\omega$ and $\lambda\rightarrow-\lambda$. Changing the sign of $\lambda$ indirectly changes the sign of $\lambda_{1}$, thus retrieving the Hamiltonian matrix in Eq. (39) of the even sector, as given in Eq. (2). With the transformed parameters, the eigenstates of Eq. (41) simplify, and we obtain the only contributing eigenvector of the form |$\ket{E^{o}_{3}}=(|001\rangle+|111\rangle)/\sqrt{2}$, which is the same as the initial state $|\psi^{o}(0)\rangle$. The remaining eigenvectors are orthogonal to the initial state and thus do not contribute to the dynamics. Furthermore, since the eigenvector $E^{o}_{3}$ has zero energy, the dynamics of the initial state in Eq. (42) do not evolve at all, according to Eq. (18). Consequently, we obtain the same results as shown in Fig. 2(a), which are also displayed in Fig. 6(b) at $\lambda=-1$ and $\omega=0$ for the odd sector.
However, in the case of asymmetric couplings ($\lambda\neq 1$), it is important to note that the coefficients $\eta^{\prime}$ and $\chi^{\prime}$ are interchanged with their corresponding configurations compared to the results in Eq. (19) for the even sector. This interchange occurs primarily to preserve the odd parity of the system. Therefore, the dynamics remain the same as shown in Fig. 6(d) at $\lambda=-0.7$ and $\omega=0$. Similar changes are expected to occur for finite $\omega$ values, which we do not show here. Consequently, we conclude that the measurable quantities exhibit different dynamics for $\omega=0$ and $\omega\neq 0$, similar to the even sector, but with differences in periodicity. More importantly, the overall characteristics of the dynamics in both sectors remain consistent and can be analyzed independently, as dictated by the odd-even symmetry of the system.